# Learning Optimal Stable Matches in Decentralized Markets with Unknown Preferences

### v1

@@ VII-C Proof
All that remains to be shown is that a minimum resistance tree rooted at any other node $z\neq e^{*}$ has resistance strictly greater than $\rho(e^{*})$, the stochastic potential of $e^{*}$. First, consider any $z\in\mathbf{C}\cup\mathbf{D}$, and let $T(z)$ be a minimum resistance tree rooted at $z$. Every vertex $z^{\prime}\in\mathbf{Z}\setminus\{z,e^{*}\}$ contributes resistance at least $r^{*}(z^{\prime})$, the vertex $e^{*}$ contributes resistance at least $2$, and the resistance corresponding to $z$ is removed, so the stochastic potential of $T(z)$ is at least $\rho(z)\geq\rho(e^{*})+2-r^{*}(z)$. Because $z$ is not an equilibrium, $r^{*}(z)$ is strictly less than $2$, so $\rho(z)>\rho(e^{*})$. Next, consider a state $e\in\mathbf{E}$. Since $r^{*}(e)=r^{*}(e^{*})=2$, we have that $\rho(e)\geq\rho(e^{*})$. Because $T(e)$ is a tree rooted at $e$, there exists a unique path from $e$ to $e^{*}$. However, Corollary 1 implies that every best response sequence starting from a near-stable match with respect to $\mu^{*}$ must lead back to $\mu^{*}$. Thus, the path from $e$ to $e^{*}$ in $T(e)$ must include at least one transition, say, from $z$ to $z^{\prime}$, which is not a best response update (if $z\in\mathbf{C}\cup\mathbf{D}$) or does not lead to a near-stable match (if $z\in\mathbf{E}$). This transition is not easy, so it has resistance strictly greater than $r^{*}(z)$; thus, $\rho(e)>\rho(e^{*})$. This establishes that the unique recurrence class of minimum stochastic potential is $e^{*}$, and hence, the POSM is the only stochastically stable state.

### v2

@@ VI Proof of Theorem
All that remains to be shown is that a minimum resistance tree rooted at any other node $z\neq e^{*}$ has resistance strictly greater than $\rho(e^{*})$, the stochastic potential of $e^{*}$. First, consider any $z\in\mathbf{C}\cup\mathbf{D}$, and let $T(z)$ be a minimum resistance tree rooted at $z$. Every vertex $z^{\prime}\in\mathbf{Z}\setminus\{z,e^{*}\}$ contributes resistance at least $r^{*}(z^{\prime})$, the vertex $e^{*}$ contributes resistance at least $2$, and the resistance corresponding to $z$ is removed, so the stochastic potential of $T(z)$ is at least $\rho(z)\geq\rho(e^{*})+2-r^{*}(z)$. Because $z$ is not an equilibrium, $r^{*}(z)$ is strictly less than $2$, so $\rho(z)>\rho(e^{*})$. Next, consider a state $e\in\mathbf{E}$. Since $r^{*}(e)=r^{*}(e^{*})=2$, we have that $\rho(e)\geq\rho(e^{*})$. Because $T(e)$ is a tree rooted at $e$, there exists a unique path from $e$ to $e^{*}$. However, [30, Corollary 1] implies that every best response sequence starting from a near-stable match with respect to $\mu^{*}$ must lead back to $\mu^{*}$. Thus, the path from $e$ to $e^{*}$ in $T(e)$ must include at least one transition, say, from $z$ to $z^{\prime}$, which is not a best response update (if $z\in\mathbf{C}\cup\mathbf{D}$) or does not lead to a near-stable match (if $z\in\mathbf{E}$). This transition is not easy, so it has resistance strictly greater than $r^{*}(z)$; thus, $\rho(e)>\rho(e^{*})$. This establishes that the unique recurrence class of minimum stochastic potential is $e^{*}$, and hence, the POSM is the only stochastically stable state.